Point charges of 0.27 \mu C and 0.86 \mu C are placed 0.45 m apart.

A) At what point along the line between them is the electric field zero? Give your answer in meters from the 0.27 \mu C charge.

B) What is the magnitude of the electric field, in newtons per coulomb, halfway between them?
Answer #1

here,

charge 1 , q1 = 0.27 uC = 0.27 * 10^-6 C

charge 2 , q2 = 0.86 uC = 0.86 * 10^-6 C

r = 0.45 m

a)

let the electric feild is zero at x distance from q1

electric feild , E = K * q1 /x^2 - K * q2 /(r - x)^2

as the electric feild is zero

0 = K * q1 /x^2 - K * q2 /(r - x)^2

0 = 0.27 /x^2 - 0.86 /(0.45 - x)^2

solving for x

x = 0.16 m

the electric feild is zero at 0.16 m from 0.27 uC

b)

the magnitude of electric feild , E = K * q2 /(r/2)^2 - K * q1 /(r/2)^2

E = 9 * 10^9 * 10^-6 /(0.45/2)^2 * ( 0.86 - 0.27)

E = 1.05 * 10^5 N/C